OpenStudy (lgbasallote):

how to find the derivative of absolute value? Can someone give a demo? (explanations appreciated)

OpenStudy (unklerhaukus):

dosent the function have to be simply connected?

OpenStudy (lgbasallote):

isnt that the integral?

OpenStudy (ash2326):

|x| is not differentiable at x=0 so we can't differentiate it

OpenStudy (lgbasallote):

then when is it differentiable?

OpenStudy (unklerhaukus):

thats what i ment to say @ash2326

OpenStudy (ash2326):

It's differentiable everywhere except x=0

OpenStudy (ash2326):

Got your point @UnkleRhaukus

OpenStudy (ash2326):

However \( x|x| \) is differentiable at x=0

OpenStudy (lgbasallote):

is it because at (0,0) it's no longer a curve but a point?

OpenStudy (lgbasallote):

oh btw @ash2326 @UnkleRhaukus i was asking how to find the derivative lol

OpenStudy (anonymous):

\[(\left| u \right|)'=\frac{u u'}{\left| u \right|}\]

OpenStudy (lgbasallote):

@mukushla so (|f(x))' = \(\frac{f(x) f'(x)}{|f(x)|}?\)

OpenStudy (anonymous):

yes do u wanna prove it?

OpenStudy (lgbasallote):

i dont think im in the mood for proving :p lol

OpenStudy (anonymous):

see this \[\frac{d}{dx} \left| u \right| = \frac{d}{dx}\sqrt{u^2} =\frac{d}{dx} (u^2)^{1/2}=\frac{1}{2} \ 2 u u' (u^2)^{-1/2}=\frac{u u'}{(u^2)^{1/2}}=\frac{u u'}{\left| u \right|}\]

OpenStudy (lgbasallote):

lol why =_=